2 + 2(y-8) = y - 4
Answer by elima(1433)

2 + 2(y-8) = y - 4
First distribute the 2;
2+2y-16=y-4
Now subtract 2-16;
2y-14=y-4
subtract both sides by y;
2y-14-y=y-4-y
y-14=-4
Add 14 to both sides;
y-14+14=-4+14
y=10
Hope you understand
=)